# Computing ternary liquid phase diagrams: Fe-Cu-Ni

### v2

@@ Abstract
Figure 4: Liquid binodal (orange) and spinodal (blue) decomposition points for Fe-Cu. The curves are computed directly from the analytic fits to free energy. The critical temperature is 2286.7K and critical Cu concentration is 33.775 at.%.
@@
Computing the phase diagram for a ternary is more complex than the binary; we map out the boundary iteratively with a predictor-corrector method at each temperature, by proceeding from the Fe-Cu binary into the ternary phase field. We take the two binodal points, and given the Gibbs free energy, chemical potential (first derivatives) and derivatives of the chemical potentials, we make a small change in each concentration that should (a) keep the chemical potentials equal to first order, and (b) maintain a common tangent to first order. After that initial step, we correct each concentration along directions that make the chemical potentials match, and push to the common tangent. This procedure identifies a tie line, and we continue stepping along until the binodal collapses at the critical point. The full algorithm is described in Appendix A. Along each tie-line, the spinodal points are identified as the points where the second derivative of $G$ is zero via numerical root-finding.
@@
Figure 8: Molecular dynamics snapshots Fe50Cu25Ni25 at 2100K (top), 1800K (middle), 1500K (bottom). In each snapshot, Fe (left), Cu (middle), and Ni (right) are shown separately.
@@
This is the $S^{0}$ method. To find $S^{0}$ from $S(\mathbf{k})$, we fit the small $\mathbf{k}$ behavior to the functional form $S^{0}/(1+k^{2}\xi^{2})$ for length scale $\xi$. We can directly compare the calculation of the second derivative of the Gibbs free energy from our fits to that determined via the $S^{0}$ method in Fig. 9. As can be seen for the three binary subsystems, the $S^{0}$ calculations of the second derivatives qualitatively agree with the analytic forms, but show significant quantitative disagreement. While the disagreement is most significant for the Fe-Cu system, especially as it undergoes partial phase separation, there are disagreements for all temperatures in the Cu-Ni and Ni-Fe systems as well. While we did not attempt to fit analytic functions to the $S^{0}$ data, it seems unlikely to reproduce the binary coefficients in Fig. 2. It is possible that this may be remedied by significantly longer molecular dynamics trajectories. In addition, as the $S^{0}$ method only produces derivatives of the chemical potentials, additional computation would be required to provide a baseline chemical potential, which is necessary to compute a phase diagram.
@@
Fig. 11 shows the absolute Gibbs free energies of the pure liquid states, calculated using non-equilibrium Hamiltonian integration. While all of the data needed to compute the liquid phase diagram is available from a virtual semigrand canonical Widom approach, computing the absolute Gibbs free energy for the full phase diagram requires additional reference data. The differences in chemical potentials allows us to determine the Gibbs free energy up to an additive constant for each temperature; in order to compute equilibria against other phases, we would need to find a reference. Appendix B gives the details of our computation of the absolute Gibbs free energy for pure liquid Fe, Cu, and Ni, performed at 3000K, 2700K, 2400K, 2100K, and 1800K. The standard error in the Gibbs free energy from integration is less than 0.6 meV for all of the simulations. Similarly, the Gibbs free energies all fit our temperature dependence with errors below 0.1 meV. When we compare the difference in absolute Gibbs free energies for our pure liquids to that given by our coefficients in Fig. 2, we find differences of 1 meV or lower. If we extrapolate the fit down to 1200K, we find a difference of 10 meV for Fe-Cu. In addition, we can take the thermal average energy as a function of temperature from our runs, and integrate to get the difference in $G(T)$ at different temperatures; compared with the non-equilibrium Hamiltonian integration, the error is 1 meV over the 1800–3000K temperature range. So while we have benchmarked our results, in a practical calculation, one would need to do a single absolute Gibbs free energy calculation at one reference temperature, and then use the energy integration and chemical potential differences to have absolute Gibbs free energies across the entire chemical and temperature space.
@@
We have demonstrated the prediction of phase separation in a ternary liquid system using a virtual semigrand canonical Widom approach, sampled with molecular dynamics. In addition the determination of the Gibbs free energy for the ternary Fe-Cu-Ni system as a function of composition and temperature over the range 1200–3000K, we mapped the binodal and spinodal limits in the Gibbs triangle. Using molecular dynamics, we could also determine density, pair correlation functions, and static structure factors. The latter was used to provide benchmark computation of the second derivative of the Gibbs free energy via the $S^{0}$ method; it appears that longer simulations may be necessary to provide sufficient accuracy to be useful for fitting of Gibbs free energies. The virtual semigrand canonical Widom approach has overhead of roughly a factor of 2 with a regular molecular dynamics run; that said, the standard error estimates shown here indicate that shorter trajectories or fewer samples could be used. In addition, a hybrid Monte Carlo / molecular dynamics approach may further increase the efficiency of sampling, albeit at the expense of full trajectory information. We also benchmarked our computation against absolute Gibbs free energy calculations for the pure liquid; while these were computationally intensive, they show that the computed free energy matches to well small fractions of the thermal energy. While the numerical results hinge on the accuracy of the underlying EAM potential to describe the interaction of Fe, Cu, and Ni, the trends identified should be useful for the design of ternary alloys in additive manufacturing. Moreover, this serves as a useful benchmark for future, more accurate, computation of phase diagrams.
@@ Acknowledgements.
Appendix A Predictor-corrector algorithm for ternary binodal points and tie-lines
@@
We begin our identification of the binodal points in the ternary from the known binodal points in the Fe-Cu binary. Initially, our two concentrations obey the common tangent construction and have equal chemical potentials only along the Fe-Cu axis, as $x_{\text{Ni}}=0$. As a first step, we introduce a small $x_{\text{Ni}}=10^{-4}$ concentration, split between the two binodal points to make the full chemical potentials equal. Then, let our initial binodal concentration vectors be $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ with Gibbs free energies $G_{1}$ and $G_{2}$, chemical potentials $\boldsymbol{\mu}_{1}$ and $\boldsymbol{\mu}_{2}$, and second derivative matrices $\underline{h}_{1}:=\boldsymbol{\nabla}\boldsymbol{\nabla}G_{1}$ and $\underline{h}_{2}:=\boldsymbol{\nabla}\boldsymbol{\nabla}G_{2}$. Since we start on the binodal, $\boldsymbol{\mu}_{1}=\boldsymbol{\mu}_{2}=\boldsymbol{\mu}$. Our predictor step corresponds with $\mathbf{x}_{1}\to\mathbf{x}_{1}+\boldsymbol{\Delta}\mathbf{x}_{1}$ and $\mathbf{x}_{2}\to\mathbf{x}_{2}+\boldsymbol{\Delta}\mathbf{x}_{2}$; maintaining equal chemical potentials requires the constraint
$\underline{h}_{1}\cdot\boldsymbol{\Delta}\mathbf{x}_{1}=\underline{h}_{2}\cdot\boldsymbol{\Delta}\mathbf{x}_{2}.$ (10)
The common tangent requirement (difference in Gibbs free energies must equal the chemical potential times the concentration difference) is
@@
$\begin{split}G_{2}+\boldsymbol{\mu}\cdot\boldsymbol{\Delta}\mathbf{x}_{2}-G_{1}-\boldsymbol{\mu}\cdot\boldsymbol{\Delta}\mathbf{x}_{1}&=(\boldsymbol{\mu}+\underline{h}_{1}\cdot\boldsymbol{\Delta}\mathbf{x}_{1})\cdot(\mathbf{x}_{2}+\boldsymbol{\Delta}\mathbf{x}_{2}-\mathbf{x}_{1}-\boldsymbol{\Delta}\mathbf{x}_{1})\\ G_{2}-G_{1}+\boldsymbol{\mu}\cdot(\boldsymbol{\Delta}\mathbf{x}_{2}-\boldsymbol{\Delta}\mathbf{x}_{1})&=\boldsymbol{\mu}\cdot(\mathbf{x}_{2}-\mathbf{x}_{1})+\boldsymbol{\mu}\cdot(\boldsymbol{\Delta}\mathbf{x}_{2}-\boldsymbol{\Delta}\mathbf{x}_{1})+(\mathbf{x}_{2}-\mathbf{x}_{1})\cdot\underline{h}_{1}\cdot\boldsymbol{\Delta}\mathbf{x}_{1}+O(\Delta x^{2})\\ 0&=(\mathbf{x}_{2}-\mathbf{x}_{1})\cdot\underline{h}_{1}\cdot\boldsymbol{\Delta}\mathbf{x}_{1}\end{split}$ (11)
and similarly $(\mathbf{x}_{2}-\mathbf{x}_{1})\cdot\underline{h}_{2}\cdot\boldsymbol{\Delta}\mathbf{x}_{2}=0$. Thus, we move $\boldsymbol{\Delta}\mathbf{x}_{1}$ in a direction perpendicular to $\underline{h}_{1}\cdot(\mathbf{x}_{2}-\mathbf{x}_{1})$ and $\boldsymbol{\Delta}\mathbf{x}_{2}$ perpendicular to $\underline{h}_{2}\cdot(\mathbf{x}_{2}-\mathbf{x}_{1})$; the magnitudes are chosen to ensure Eqn. 10 holds and produce a total step size given by $\Delta x_{1}^{2}+\Delta x_{2}^{2}=\Delta^{2}$. After the predictor step, we recompute $G_{1}$, $G_{2}$, $\boldsymbol{\mu}_{1}$, $\boldsymbol{\mu}_{2}$, $\underline{h}_{1}$ and $\underline{h}_{2}$; our chemical potentials are only equal to first order in our step size. Define $\Delta G:=G_{2}-G_{1}$ and $\boldsymbol{\Delta}\boldsymbol{\mu}:=\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1}$; we seek new changes in concentration $\boldsymbol{\Delta}\mathbf{x}_{1}$ and $\boldsymbol{\Delta}\mathbf{x}_{2}$ to make our chemical potentials equal and maintain the common tangent construction. We have
@@
$\begin{split}\boldsymbol{\mu}_{1}+\underline{h}_{1}\cdot\boldsymbol{\Delta}\mathbf{x}_{1}&=\boldsymbol{\mu}_{2}+\underline{h}_{2}\cdot\boldsymbol{\Delta}\mathbf{x}_{2}\\ \underline{h}_{1}\cdot\boldsymbol{\Delta}\mathbf{x}_{1}-\underline{h}_{2}\cdot\boldsymbol{\Delta}\mathbf{x}_{2}&=\boldsymbol{\Delta}\boldsymbol{\mu}\end{split}$ (12)
which gives one condition to be satisfied. Let $\mathbf{d}_{1}:=(\underline{h}_{1})^{-1}\cdot\boldsymbol{\Delta}\boldsymbol{\mu}$ and $\mathbf{d}_{2}:=-(\underline{h}_{2})^{-1}\cdot\boldsymbol{\Delta}\boldsymbol{\mu}$; if our concentration changes are $\boldsymbol{\Delta}\mathbf{x}_{1}=\alpha\mathbf{d}_{1}$ and $\boldsymbol{\Delta}\mathbf{x}_{2}=(1-\alpha)\mathbf{d}_{2}$, we will satisfy Eqn. 12 for any value of $\alpha$. The corrected chemical potential $\boldsymbol{\mu}=\boldsymbol{\mu}_{1}+\alpha\boldsymbol{\Delta}\boldsymbol{\mu}=\alpha\boldsymbol{\mu}_{2}+(1-\alpha)\boldsymbol{\mu}_{1}$. The common tangent condition is, to second order in $\mathbf{d}_{1}$ and $\mathbf{d}_{2}$,
$G_{2}+(1-\alpha)\boldsymbol{\mu}_{2}\cdot\mathbf{d}_{2}+\frac{1}{2}(1-\alpha)^{2}\mathbf{d}_{2}\cdot\underline{h}_{2}\cdot\mathbf{d}_{2}-G_{1}-\alpha\boldsymbol{\mu}_{1}\cdot\mathbf{d}_{1}+\frac{1}{2}\alpha^{2}\mathbf{d}_{1}\cdot\underline{h}_{1}\cdot\mathbf{d}_{1}=\left(\mathbf{x}_{2}+(1-\alpha)\mathbf{d}_{2}-\mathbf{x}_{1}-\alpha\mathbf{d}_{1}\right)\cdot(\boldsymbol{\mu}_{1}+\alpha\boldsymbol{\Delta}\boldsymbol{\mu}_{1})$
This can be simplified by defining $M_{1}:=\mathbf{d}_{1}\cdot\underline{h}_{1}\cdot\mathbf{d}_{1}=\boldsymbol{\Delta}\boldsymbol{\mu}\cdot\mathbf{d}_{1}$ and $M_{2}:=\mathbf{d}_{2}\cdot\underline{h}_{2}\cdot\mathbf{d}_{2}=-\boldsymbol{\Delta}\boldsymbol{\mu}\cdot\mathbf{d}_{2}$. Then, following some algebra,
$\Delta G-(\mathbf{x}_{2}-\mathbf{x}_{1})\cdot\boldsymbol{\mu}_{1}=\frac{1}{2}(1-\alpha)^{2}M_{2}-\frac{1}{2}\alpha^{2}M_{1}+\alpha(\mathbf{x}_{2}-\mathbf{x}_{1})\cdot\boldsymbol{\Delta}\boldsymbol{\mu}$ (13)
As the $\alpha^{2}$ term is $\frac{1}{2}(M_{2}-M_{1})$, this quadratic equation is most stably solved using the quadratic formula with the square root in the denominator; the sign is chosen to match the sign of the linear term in $\alpha$. With a few iterations, Eqn. 13 converges, and the tie-line is determined. The algorithm continues until the binodal points converge on each other near the critical point.

### v3

@@ III.1 Chemical potentials
As a benchmark test of our chemical potential values for the ternary compositions, we compare against semi-grand canonical Monte Carlo simulations[30]. Taking the differences in excess chemical potentials extracted from our virtual semi-grand canonical Widom calculations at the ternary compositions, we performed hybrid molecular dynamics / semi-grand canonical MC simulations in LAMMPS, where every 2048 steps, 2048 trial moves are attempted to swap atomic identities. After $2^{17}$ fs, the chemical composition is averaged, and compared with the nominal composition. For all temperatures above 2100K, the disagreement is less than 5%; starting at 2100K and below, there is disagreement in compositions in the two-phase regime, as the SGCMC method will favor one side of the two-phase separation.
@@
III.2 Phase diagrams
Figure 4: Liquid binodal (orange) and spinodal (blue) decomposition points for Fe-Cu, compared with experiment (dashed). The curves are computed directly from the analytic fits to free energy. The EAM potential critical temperature is 2286.7K and critical Cu concentration is 33.775 at.%; the experimental curves come from CALPHAD[7] where the temperature has been scaled to match the EAM critical temperature.
@@ III.2 Phase diagrams
Computing the phase diagram for a ternary is more complex than the binary; we map out the boundary iteratively at each temperature, by proceeding from the Fe-Cu binary into the ternary phase field. We take the two binodal points, and given the Gibbs free energy, chemical potential (first derivatives) and derivatives of the chemical potentials, we make a small change in each concentration to (a) keep the chemical potentials equal to first order, and (b) maintain a common tangent to first order. This requires solving a coupled quadratic equation. This procedure identifies a tie line, and we continue stepping along until the binodal collapses at the critical point. The full algorithm is described in Appendix A; alternate algorithms to map the phase boundaries from a parameterized free-energy function are available in CALPHAD software such as OpenCalphad[31]. Along each tie-line, the spinodal points are identified as the points where the second derivative of $G$ is zero via numerical root-finding.
@@ III.3 Structural analysis
Figure 8: Molecular dynamics snapshots Fe50Cu25Ni25 at 2100K (left), 1800K (middle), 1500K (right). For clarity, a slice of thickness 7.8Å (corresponding to the dip in $g(r)$ between 3rd and 4th neighbor peaks) is shown with Fe (red), Cu (copper), and Ni (green).
@@ III.4 $S^{0}$ method
This is the $S^{0}$ method. To find $S^{0}$ from $S(\mathbf{k})$, we fit the small $\mathbf{k}$ behavior to the functional form $S^{0}/(1+k^{2}\xi^{2})$ for length scale $\xi$. We can directly compare the calculation of the second derivative of the Gibbs free energy from our fits to that determined via the $S^{0}$ method in Fig. 10. As can be seen for the three binary subsystems, the $S^{0}$ calculations of the second derivatives qualitatively agree with the analytic forms, but show significant quantitative disagreement. While the disagreement is most significant for the Fe-Cu system, especially as it undergoes partial phase separation where $S^{0}$ will predict $\partial^{2}G/\partial x^{2}=0$, there are disagreements for all temperatures in the Cu-Ni and Ni-Fe systems as well. While we did not attempt to fit analytic functions to the $S^{0}$ data, it seems unlikely to reproduce the binary coefficients in Fig. 2. It is possible that this may be remedied by significantly longer molecular dynamics trajectories and larger cells for a more accurate calculation of the small $k$-limit. In addition, as the $S^{0}$ method only produces derivatives of the chemical potentials, additional computation would be required to provide a baseline chemical potential, which is necessary to compute a phase diagram.
III.5 Absolute Gibbs free energy
@@
Fig. 11 shows the absolute Gibbs free energies of the pure liquid states, calculated using non-equilibrium Hamiltonian integration. While all of the data needed to compute the liquid phase diagram is available from a virtual semigrand canonical Widom approach, computing the absolute Gibbs free energy for the full phase diagram requires additional reference data. The differences in chemical potentials allows us to determine the Gibbs free energy up to an additive constant for each temperature; in order to compute equilibria against other phases, we would need to find a reference. Appendix B gives the details of our computation of the absolute Gibbs free energy for pure liquid Fe, Cu, and Ni, performed at 3000K, 2700K, 2400K, 2100K, and 1800K. The standard error in the Gibbs free energy from integration is less than 0.6 meV for all of the simulations. Similarly, the Gibbs free energies all fit our temperature dependence with errors below 0.1 meV. When we compare the difference in absolute Gibbs free energies for our pure liquids to that given by our coefficients in Fig. 2, we find differences of 1 meV or lower. If we extrapolate the fit down to 1200K, we find a difference of 10 meV for Fe-Cu. In addition, we can take the thermal average energy as a function of temperature from our runs, and integrate to get the difference in $G(T)$ at different temperatures; compared with the non-equilibrium Hamiltonian integration, the error is 1 meV over the 1800–3000K temperature range. So while we have benchmarked our results, in a practical calculation, one would need to do a single absolute Gibbs free energy calculation at one reference temperature, and then use the energy integration and chemical potential differences to have absolute Gibbs free energies across the entire chemical and temperature space. Note also that each non-equilibrium Hamiltonian integration calculation required 2 million energy evaluations for each composition and temperature, compared with approximately 260,000 energy evaluations for the virtual semigrand canonical Widom method including the trajectory and the calculation of energy changes.
IV Conclusions
@@
We have demonstrated the prediction of phase separation in a ternary liquid system using a virtual semigrand canonical Widom approach, sampled with molecular dynamics. In addition the determination of the Gibbs free energy for the ternary Fe-Cu-Ni system as a function of composition and temperature over the range 1200–3000K, we mapped the binodal and spinodal limits in the Gibbs triangle. Using molecular dynamics, we could also determine density, pair correlation functions, and static structure factors. The latter was used to provide benchmark computation of the second derivative of the Gibbs free energy via the $S^{0}$ method; it appears that longer simulations with larger cells may be necessary to provide sufficient accuracy to be useful for fitting of Gibbs free energies. The virtual semigrand canonical Widom approach has overhead of roughly a factor of 2 with a regular molecular dynamics run; that said, the standard error estimates shown here indicate that shorter trajectories or fewer samples could be used. In addition, a hybrid Monte Carlo / molecular dynamics approach may further increase the efficiency of sampling, albeit at the expense of full trajectory information. We also benchmarked our computation against absolute Gibbs free energy calculations for the pure liquid; while these were computationally intensive, they show that the computed free energy matches to well small fractions of the thermal energy. The method is not restricted to liquids, and can be applied to multicomponent solid systems as well. While the numerical results hinge on the accuracy of the underlying EAM potential to describe the interaction of Fe, Cu, and Ni, the trends identified should be useful for the design of ternary alloys in additive manufacturing. Moreover, this serves as a useful benchmark for future, more accurate, computation of phase diagrams.
@@ Acknowledgements.
Appendix A Iterative algorithm for ternary binodal points and tie-lines
@@
We begin our identification of the binodal points in the ternary from the known binodal points in the Fe-Cu binary. Initially, our two concentrations obey the common tangent construction and have equal chemical potentials only along the Fe-Cu axis, as $x_{\text{Ni}}=0$. As a first step, we introduce a small $x_{\text{Ni}}=10^{-4}$ concentration, but split between the two binodal points to make the full chemical potentials equal. Then, let our initial binodal concentration vectors be $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ with Gibbs free energies $G_{1}$ and $G_{2}$, chemical potentials $\boldsymbol{\mu}_{1}$ and $\boldsymbol{\mu}_{2}$, and second derivative matrices $\underline{h}_{1}:=\boldsymbol{\nabla}\boldsymbol{\nabla}G_{1}$ and $\underline{h}_{2}:=\boldsymbol{\nabla}\boldsymbol{\nabla}G_{2}$. We define $\boldsymbol{\Delta}\mathbf{x}:=\mathbf{x}_{2}-\mathbf{x}_{1}$ as the tie-line direction, and identify two sets of (normalized) directions: $\hat{\mathbf{p}}_{i}:=(\underline{h}_{i}\cdot\boldsymbol{\Delta}\mathbf{x})/|\underline{h}_{i}\cdot\boldsymbol{\Delta}\mathbf{x}|$ and $\hat{\mathbf{t}}_{i}$ that is perpendicular to $\hat{\mathbf{p}}_{i}$ in the Gibbs triangle plane. We change our compositions with $\mathbf{x}_{1}\to\mathbf{x}_{1}+(s-\alpha)\hat{\mathbf{t}}_{1}+\beta_{1}\hat{\mathbf{p}}_{1}$ and $\mathbf{x}_{2}\to\mathbf{x}_{2}+(s+\alpha)\hat{\mathbf{t}}_{2}+\beta_{2}\hat{\mathbf{p}}_{2}$ with a step size $s$. The three unknowns, $\alpha$, $\beta_{1}$, and $\beta_{2}$ are chosen to guarantee that the chemical potentials are equal to first order and that the new points lie on the common tangent to second order. The chemical potential difference is
$0=\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1}+\underline{h}_{2}\cdot((s+\alpha)\hat{\mathbf{t}}_{2}+\beta_{2}\hat{\mathbf{p}}_{2})-\underline{h}_{1}\cdot((s-\alpha)\hat{\mathbf{t}}_{1}+\beta_{1}\hat{\mathbf{p}}_{1})$ (10)
whose equality we check by dotting into $\boldsymbol{\Delta}\mathbf{x}$ and a direction $\mathbf{n}$ that is perpendicular in the Gibbs triangle plane. Defining $\boldsymbol{\Delta}\boldsymbol{\mu}:=\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1}$ (the small error in chemical-potential equality before changing composition), we have
@@
$\begin{split}\boldsymbol{\Delta}\mathbf{x}\cdot\boldsymbol{\Delta}\boldsymbol{\mu}=&\ (\underline{h}_{1}\boldsymbol{\Delta}\mathbf{x})\cdot((s-\alpha)\hat{\mathbf{t}}_{1}+\beta_{1}\hat{\mathbf{p}}_{1})\\ &-(\underline{h}_{2}\boldsymbol{\Delta}\mathbf{x})\cdot((s+\alpha)\hat{\mathbf{t}}_{2}+\beta_{2}\hat{\mathbf{p}}_{2})\\ =&\ (\boldsymbol{\Delta}\mathbf{x}\underline{h}_{1}\hat{\mathbf{p}}_{1})\beta_{1}-(\boldsymbol{\Delta}\mathbf{x}\underline{h}_{2}\hat{\mathbf{p}}_{2})\beta_{2}\end{split}$ (11)
and
@@
$\begin{split}\mathbf{n}\cdot\boldsymbol{\Delta}\boldsymbol{\mu}=&\ (\underline{h}_{1}\mathbf{n})\cdot((s-\alpha)\hat{\mathbf{t}}_{1}+\beta_{1}\hat{\mathbf{p}}_{1})\\ &-(\underline{h}_{2}\mathbf{n})\cdot((s+\alpha)\hat{\mathbf{t}}_{2}+\beta_{2}\hat{\mathbf{p}}_{2})\\ =&\ (\mathbf{n}\underline{h}_{1}\hat{\mathbf{t}}_{1}-\mathbf{n}\underline{h}_{2}\hat{\mathbf{t}}_{2})s-(\mathbf{n}\underline{h}_{1}\hat{\mathbf{t}}_{1}+\mathbf{n}\underline{h}_{2}\hat{\mathbf{t}}_{2})\alpha\\ &+(\mathbf{n}\underline{h}_{1}\hat{\mathbf{p}}_{1})\beta_{1}-(\mathbf{n}\underline{h}_{2}\hat{\mathbf{p}}_{2})\beta_{2}\end{split}$ (12)
These two equations are linear in our three unknowns.
The final equation is the common-tangent requirement. We define $\Delta G_{i}:=G_{2}-G_{1}-\boldsymbol{\mu}_{i}\cdot\boldsymbol{\Delta}\mathbf{x}$, which is the small error in the common-tangent before changing composition. The common-tangent equation is written most compactly with the composition changes $\boldsymbol{\Delta}\mathbf{x}_{i}:=(s\mp\alpha)\hat{\mathbf{t}}_{i}+\beta_{i}\hat{\mathbf{p}}_{i}$ as
$G_{2}+\boldsymbol{\mu}_{2}\cdot\boldsymbol{\Delta}\mathbf{x}_{2}+\frac{1}{2}\boldsymbol{\Delta}\mathbf{x}_{2}\underline{h}_{2}\boldsymbol{\Delta}\mathbf{x}_{2}-G_{1}-\boldsymbol{\mu}_{1}\cdot\boldsymbol{\Delta}\mathbf{x}_{1}-\frac{1}{2}\boldsymbol{\Delta}\mathbf{x}_{1}\underline{h}_{1}\boldsymbol{\Delta}\mathbf{x}_{1}=(\boldsymbol{\mu}_{i}+\underline{h}_{i}\boldsymbol{\Delta}\mathbf{x}_{i})\cdot(\mathbf{x}_{2}-\mathbf{x}_{1}+\boldsymbol{\Delta}\mathbf{x}_{2}-\boldsymbol{\Delta}\mathbf{x}_{1})$ (13)
which holds for $i=1,2$. With some algebraic manipulation and the requirement that $\boldsymbol{\mu}_{1}+\underline{h}_{1}\boldsymbol{\Delta}\mathbf{x}_{1}=\boldsymbol{\mu}_{2}+\underline{h}_{2}\boldsymbol{\Delta}\mathbf{x}_{2}$, Eqn. 13 becomes
$\Delta G_{i}-\boldsymbol{\Delta}\mathbf{x}\cdot\underline{h}_{i}\boldsymbol{\Delta}\mathbf{x}_{i}=\frac{1}{2}\boldsymbol{\Delta}\mathbf{x}_{2}\underline{h}_{2}\boldsymbol{\Delta}\mathbf{x}_{2}-\frac{1}{2}\boldsymbol{\Delta}\mathbf{x}_{1}\underline{h}_{1}\boldsymbol{\Delta}\mathbf{x}_{1}$ (14)
which can be summed over $i=1,2$ to give
$\Delta G_{1}+\Delta G_{2}=\boldsymbol{\Delta}\mathbf{x}\cdot\underline{h}_{1}\boldsymbol{\Delta}\mathbf{x}_{1}+\boldsymbol{\Delta}\mathbf{x}\cdot\underline{h}_{2}\boldsymbol{\Delta}\mathbf{x}_{2}+\boldsymbol{\Delta}\mathbf{x}_{2}\underline{h}_{2}\boldsymbol{\Delta}\mathbf{x}_{2}-\boldsymbol{\Delta}\mathbf{x}_{1}\underline{h}_{1}\boldsymbol{\Delta}\mathbf{x}_{1}.$ (15)
Expanding this equation in terms of $\alpha$ and $\beta_{i}$ gives the final equation
$(2\hat{\mathbf{t}}_{1}\underline{h}_{1}\hat{\mathbf{t}}_{1}+2\hat{\mathbf{t}}_{2}\underline{h}_{2}\hat{\mathbf{t}}_{2})s\alpha+(-2\hat{\mathbf{p}}_{1}\underline{h}_{1}\hat{\mathbf{t}}_{1}s+\boldsymbol{\Delta}\mathbf{x}\underline{h}_{1}\hat{\mathbf{p}}_{1})\beta_{1}+(2\hat{\mathbf{p}}_{2}\underline{h}_{2}\hat{\mathbf{t}}_{2}s+\boldsymbol{\Delta}\mathbf{x}\underline{h}_{2}\hat{\mathbf{p}}_{2})\beta_{2}=\Delta G_{1}+\Delta G_{2}\\ +(\hat{\mathbf{t}}_{1}\underline{h}_{1}\hat{\mathbf{t}}_{1}-\hat{\mathbf{t}}_{2}\underline{h}_{2}\hat{\mathbf{t}}_{2})(s^{2}+\alpha^{2})-2\hat{\mathbf{p}}_{1}\underline{h}_{1}\hat{\mathbf{t}}_{1}\beta_{1}\alpha-2\hat{\mathbf{p}}_{2}\underline{h}_{2}\hat{\mathbf{t}}_{2}\beta_{2}\alpha+\hat{\mathbf{p}}_{1}\underline{h}_{1}\hat{\mathbf{p}}_{1}\beta_{1}^{2}-\hat{\mathbf{p}}_{2}\underline{h}_{2}\hat{\mathbf{p}}_{2}\beta_{2}^{2}$ (16)
The three equations, Eqn. 11, Eqn. 12, and Eqn. 16 can be written as a linear system in our variables $\alpha$ and $\beta_{i}$ equal to constants plus quadratic terms from Eqn. 16. This system can be solved numerically via iteration to self-consistency. With a few iterations, the solution converges, and the tie-line is determined. Because the deviations from the common-tangent and equal chemical potentials are included in the next iteration, the algorithm shows numerical stability for small values of $s$. The algorithm continues until the binodal points converge on each other near the critical point.